OpenStudy (anonymous):

what equation of the line passes through the point (-1, 7) and perpendicular to y = -1/4x - 5

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

look for a similar slope...?

OpenStudy (luigi0210):

Well here. Do you know what the perpendicular slope of \[y=-\frac{1}{4}-5\]

OpenStudy (luigi0210):

*\[y=\frac{-1}{4}x-5\]

OpenStudy (anonymous):

the slope would be -1\[\frac{ -1 }{ 4 }\]

OpenStudy (luigi0210):

Not exactly. When finding the perpendicular slope you have to take the negative reciprocal. For example the perpendicular slope of 2 would be -1/2

OpenStudy (anonymous):

so this would be 4? or 4/1?

OpenStudy (luigi0210):

Yup! :) 4/1 is the same as 4

OpenStudy (luigi0210):

So you could just use 4

OpenStudy (anonymous):

so how do I determine b?

OpenStudy (luigi0210):

Now that we have the slope we can find the equation. Do you know the point-slope formula?

OpenStudy (anonymous):

y-y1= m(x-x1)

OpenStudy (luigi0210):

Correct, and you re given a set of points correct?

OpenStudy (anonymous):

(-1,7)

OpenStudy (luigi0210):

So now we have m=4, x1=-1, and y1=7. Guess what we do next?

OpenStudy (anonymous):

plug it into the formula

OpenStudy (luigi0210):

Yup!

OpenStudy (luigi0210):

Tell me what you come up with.

OpenStudy (anonymous):

y = 4x + 3

OpenStudy (luigi0210):

Good job! See it wasn't so hard now was it? :)

OpenStudy (anonymous):

no. thanks (:

OpenStudy (luigi0210):

Yup, anytime!